OpenStudy (anonymous):

Find the volume of the solid of revolution generated by revolving the equation \(y=\sqrt(x)\) about the line x=4 (The region is bounded by the x axis and x=4)

OpenStudy (inkyvoyd):

I understand if I rewrite the equation as y=\(\sqrt{x-4}\) then integrate with appropriate bounds it will work, but there is some sort of alternate interpretation I am not understanding.

OpenStudy (inkyvoyd):

*rewrite the equation as y=\(\sqrt{x+4}\)

OpenStudy (amistre64):

 \[\int 2\pi ~r~h\]